OpenStudy (anonymous):

Given the equation y - 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form. -three fourthsx + y = 10 3x - 4y = -40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (anonymous):

@gorv @iGreen

OpenStudy (gorv):

y=mx+c

OpenStudy (gorv):

three forth=3/4

OpenStudy (gorv):

y-4= three forth x +threeforth *8

OpenStudy (gorv):

\[threeforth*8=\frac{ 3 }{ 4 }*8=6\]

OpenStudy (gorv):

y-4=threeforth x+6 add 4 both side y=threeforth x +6+4 y=threeforth x +10